# Galley: Modern Query Optimization for Sparse Tensor Programs

### v3

@@ 5.4. Identifying Minimal Sub-Queries
Fig. 5 shows the annotated expression tree for our logistic regression example after distributing the multiplication as in Eq. (2). The sigmoid function is the outermost layer of the expression, so it appears at the top of the tree. The summations all occur just inside the sigmoid function, so they annotate the addition operator, as denoted by the red arrow. Then, each multiplication has three children which are each indexed input tensors.
@@
Given the AET, we can identify the minimal sub-queries for a particular index by starting at the node where it’s annotated and traversing downwards according to the algebraic properties of each internal node. We now describe the traversal rules for functions which are distributive, non-distributive, and commutative with respect to the aggregation operator.
@@
Distributive Functions When a function which distributes over the aggregate is reached (e.g. multiplication which distributes over summation), we examine how many of the children contain the current index. Each child is a subtree of the AET, and it contains the index iff it contains a tensor that is indexed by it. If one child contains the index, we simply traverse down that child’s branch, i.e. we factor the other children out of the aggregate. When multiple children contain the index, we can’t go any further down the tree, so we wrap the sub-tree rooted at that node in an aggregate and return it as our MSQ. If the function is commutative and associative, we are slightly more precise and only include the children which contain the index.
@@
Commutative, Identical Functions When the node’s function is the same as the aggregate function and is commutative, we can push the aggregate down to each child independently. For example, if we have the expression $\sum_{i}A_{i}+B_{i}$, we can transform it into $\sum_{i}A_{i}+\sum_{i}B_{i}$. For all children which contain the index, we add the result of traversing down its branch to the MSQ and replace it with an alias to the result. If a child doesn’t contain the index, then we need to account for the repeated application of the aggregate function. To do this, we identify the function $g(x,n)=f(x,...,x)$ which represents the repeated application of our aggregate function $f$. When $f$ is addition, $g$ is multiplication. When $f$ is idempotent, $g(x,n)=x$. We then wrap each non-index child in a Map with function $g$ and the size of the index’s dimension as the second child.
@@
Blocking Functions A function which doesn’t distribute or commute with our aggregate function is called a blocking function. When we reach a blocking function in our traversal, we simply wrap it in our aggregate and return the sub-tree as our MSQ. For example, the expression $\sum_{j}\sqrt{A_{ij}B_{jk}}$ cannot be rewritten as $\sqrt{\sum_{j}A_{ij}\sum_{j}B_{jk}}$ because $\sqrt{}$ is a blocking function.
@@
Discussion Galley builds on and extends the FAQ framework for optimizing conjunctive queries with aggregation(Khamis et al., 2016). The FAQ framework explored the optimization of queries with the following form where each $\bigoplus^{(i)}$ is either equal to or forms a semi-ring with $\otimes$,
@@
While this captures many important problems, it lacks the flexibility needed to support a general tensor processing system. For example, consider a slightly modified version of the SDDMM kernel,
$\sum_{j}A_{ik}(B_{ij}+C+{jk})$
This expression is not an FAQ query because it mixes addition and multiplication within the pointwise expression. Similarly, our logistic regression example, $\sigma(\sum_{j}X_{ij}\theta_{j})$ cannot be expressed as an FAQ because the aggregate occurs within the $\sigma$. Galley’s logical optimizer extends this framework by accommodating arbitrary pointwise function composition and arbitrary placement of aggregates within expressions.
@@ 5.5. Restricted Elimination Orders
Lastly, before we can search for the optimal elimination order, we need to define the space of valid elimination orders. Depending on the structure of the input, the order in which indices can be eliminated might be restricted. This is due to two issues: (1) non-commutative aggregates and (2) aggregate placement. The former is when an aggregate wraps another aggregate which it doesn’t commute with. For example, if the expression is $\max_{i}\sum_{j}A_{ij}$, we have to perform the summation before handling the maximum because max and $\sum$ do not commute. The latter issue arises when an aggregate wraps another aggregate but cannot reach it via the traversal described above, e.g. $\sum_{i}\sqrt{\sum_{j}A_{ij}}$. In this case, the inner aggregate must be performed first. Collectively, these restrictions form a partial ordering on the index variables which needs to be respected when we enumerate elimination orders.
5.6. Search Algorithm
@@
We have now simplified the complicated issue of high level optimization to the discrete problem of choosing an optimal order on the aggregated index variables. So, in this section, we present two algorithms for searching for that optimal order using the tools described above.
@@
Greedy The greedy approach simply chooses the cheapest index to aggregate at each point. It does this by finding the minimal sub-query for each index and computing its cost. For the cheapest index, it’s minimal sub-query is removed from the expression and appended to the logical plan, then it is replaced in the remaining query with an alias to the result. This continues until no aggregates remain in the expression.
@@
Branch-and-Bound The branch-and-bound approach computes the optimal variable order and is broken into two steps. The first step uses the greedy algorithm to produce an upper bound on the cost of the overall plan. The second step performs a dynamic programming algorithm. In this algorithm, the keys of the DP table are unordered sets of indices, and the values are a tuple of ordered lists of indices, remaining queries, and costs. We initialize the table with the empty set and a cost of zero. At each step, we iterate through the entries of the DP table and attempt to aggregate out another index. We use the cost bound from the first step to prune entries from the DP table whose cost is greater than the bound which is valid because costs monotonically increase as more indices are added to the set. At the end, we return the index order associated with the full set of indices.
@@ 6. Physical Optimizer
Each query in the logical dialect roughly corresponds to a single loop nest and materialized intermediate. However, there are still several decisions which need to be made about the way the kernel is computed; (1) the loop order over the indices, (2) the format (i.e. layout) of the result, and (3) the protocol (i.e. search algorithm) for accessing each index of each input. The physical optimizer makes these decisions.
@@ 6.1. Loop Order
The loop order of a tensor kernel determines which order the inputs are accessed in. A good order will result in early pruning of iterations due to early intersection of sparse inputs. Intuitively, this is similar to selecting a variable order for a worst-case optimal join algorithm.
@@
The cost of a loop order is equal to how many iterations each level of the loop nest incurs.
@@
Consider matrix chain multiplication over three sparse matrices, $A$, $B$, and $C$.
@@
Suppose that $A$ has only a single non-zero entry and that $B$ and $C$ have 5 non-zero entries per column and per row. Then, the loop order $ijkl$ is significantly more efficient than $lkji$. In the former, the first two loops, over $i$ and $j$, only incur a single iteration because they are bounded by the size of $A$. The third and fourth occur $5$ and $5^{2}$ times respectively because there are only $5$ non-zero $k$’s per $j$ in $B$ and $5$ non-zero $l$’s per $k$ in $C$. In the latter, the first two loops iterate over the full matrix $C$ despite most of those iterations not leading to useful computation.
The last piece of our cost model is the cost of transposition. If an input’s index order is not concordant with the loop order, it must be transposed before the query can be executed. This imposes a cost that is linear in that input’s size.
Optimization Algorithm
To optimize the loop order, we combine this cost model with a branch-and-bound, dynamic programming algorithm. In the first pass, it selects the cheapest loop index at each step until reaching a full loop order. This produces an upper bound on the optimal execution cost which we use to prune loop orders in the second step. This step applies a dynamic programming algorithm. Taking inspiration from Selinger’s algorithm for join ordering, each key in the DP table is a set of index variables and a set of inputs. The former represent the loops that have been iterated so far, and the latter represents the inputs which will need to be transposed based on the optimal prefix.
@@ 6.2. Intermediate Formats
Once the loop order has been determined, we select the optimal format for each query’s output. Because we operate within the fibertree abstraction, this means that we select a format for each index of the output (e.g. dense vector, hash table, etc.). There are two factors which affect this decision: (1) the kind of writes being performed (sequential vs random) (2) the sparsity of the resulting tensor. The former is important because many formats (e.g. sorted list formats) only allow for sequential construction. This means that they can only be applied if the indices of the output form a prefix of the loop order. The latter factor balances the fact that denser formats tend to be more efficient due to their memory locality and simplicity while sparser formats are asymptotically better when dealing with significantly sparse outputs. To describe this trade-off, we hand selected sparsity cutoffs between fully sparse, bytemap, and fully dense formats. To determine a particular output index’s format, we first determine the sparsity at this index level and use our cutoffs to determine which category of formats to consider. Then, we check whether we are performing sequential or random writes and select the most efficient format that supports our write pattern.
@@ 6.3. Merge Algorithms
The final decision for the physical optimizer to make is the algorithm it will use to perform each loop’s intersection. While there are some more complex strategies, we adopt a minimal approach and select a single input to iterate over for each loop. The remaining inputs are then probed into. We make this selection by estimating the number of non-zero indices that each input has, conditioned on the indices in the outer loops. This is similar to the approach taken in (Wang et al., 2024) for optimizing WCOJ.
@@ 7. Sparsity Estimation
In this section, we describe how our system performs the sparsity estimation that guides our logical and physical optimizers. We start by discussing the subtle correspondence between sparsity estimation and cardinality estimation. Then, we describe a minimal interface for sparsity estimation inspired by this correspondence. Lastly, we discuss two implementation of this framework; the uniform estimator and the chain bound.
7.1. Sparsity & Cardinality Estimation
Sparsity estimation is highly related to cardinality estimation in databases. However, translating methods for the latter to the former requires analyzing the algebraic properties of our tensor programs. For example, let $A_{ij}$ and $B_{jk}$ be sparse matrices with a fill value of 00 and $R_{A}(I,J)$ and $R_{B}(J,K)$ be relations which store the indices of their non-zero entries. Suppose we’re performing matrix multiplication,
@@ 7.1. Sparsity & Cardinality Estimation
Then, the number of non-zero values in $C$ is precisely equal to the size of the following conjunctive query,
@@
The correspondence is due to the fact that 00 is the annhilator of multiplication (i.e. $x*0=0\forall x$), so any non-zero entry $ijk$ in the output must correspond to a non-zero $ij$ in $A$ and a non-zero $jk$ in $B$. Now, consider the following instead,
@@
In this case, a nonzero $ijk$ in the output can result from a non-zero $ij$ in $A$ or a non-zero $jk$ in $B$. In traditional relational algebra where relations are over infinite domains, this kind of disjunction would result in an infinite relation. However, tensors have finite dimensions, so we can introduce relations that represent the finite domains of each index, e.g. $D_{i}=\{1,...,n_{i}\}$. This allows us to represent the index relation of the output as,
@@
Lastly, we can translate aggregations to the tensor setting as projection operations. If we have the following statement,
@@
We can express the non-zeros entries of $C$ as,
@@ 7.2. The Sparsity Statistics Interface
We use our statistics interface to annotate an expression with stats objects at every node of the AST in a bottom-up fashion. Each stats object then represents the sparsity patterns of the intermediate tensor output from that node. Surprisingly, to support sparsity estimation over the varied workloads and arbitrary operators of tensor algebra, we only need to implement a few core functions; 1) A constructor which produces statistics from a materialized tensor. This produces statistics for Input and Alias nodes. 2) A function for annihilating Map nodes (i.e. those whose children’s fill values are the annihilator of its pointwise function) which merges the children’s statistics. 3) A function for non-annihilating Map nodes which merges the children’s statistics. 4) A function for Agg which adjusts the input’s statistics to reflect an aggregation over some set of indices. 5) An estimation procedure that estimates the number of non-fill entries based on statistics about a tensor.

### v4

@@ 5.4. Identifying Minimal Sub-Expressions
Fig. 5 shows the annotated expression tree for our logistic regression example after distributing the multiplication as in Eq. (2). The sigmoid function is the outermost layer of the expression, so it appears at the top of the tree. The summations all occur just inside the sigmoid function, so they annotate the top multiplication operator, as denoted by the red arrow.
@@
Given the AET, Galley identifies an index’s MSEs by starting at the node where it is annotated and traversing downwards according to the algebraic properties of each internal node. We now describe the traversal rules for functions that are distributive, non-distributive, and commutative with respect to the aggregation operator.
@@
Distributive Functions. When we reach a function that distributes over the aggregate (e.g., * and $\sum$), we examine how many of the children, subtrees of the AET, contain the current index. If one child contains the index, we traverse down that child’s branch, i.e., we factor the other children out of the aggregate. If multiple children contain the index, we cannot go any further down the tree, so we wrap the sub-tree rooted at that node in the aggregate and return it as our MSE. If the function is commutative and associative, we are slightly more precise and include only the children that contain the index.
@@
Commutative, Identical Functions. When the node’s function is the same as the aggregate function and is commutative, we can push the aggregate down to each child independently. For example, we can transform the expression $\sum_{i}A_{i}+B_{i}$ into $\sum_{i}A_{i}+\sum_{i}B_{i}$. For all children that contain the index, we add the result of traversing down its branch to the list of MSEs and replace it with an alias to the result. If a child does not contain the index, then we need to account for the repeated application of the aggregate function. To do this, we require a function $g(x,n)=f(x,...,x)$ that represents the repeated application of our aggregate function $f$. When $f$ is addition, $g$ is multiplication. When $f$ is idempotent, $g(x,n)=x$. We then wrap each non-index child in a Map with function $g$ and the size of the index’s dimension as the second child. Further, we can choose to push down along some of these dimensions and not others, e.g. $\sum_{i}(A_{i}+B_{i}+C_{i})=\sum_{i}(A_{i}+B_{i})+\sum_{i}(C_{i})$. To allow this, we represent each path that a variable can be pushed down with a virtual index, and we include these in our elimination ordering. This increases our optimization space but allows us to capture more aggregation strategies.
@@
Blocking Functions. A function that does not distribute or commute with our aggregate function is called a blocking function. When we reach a blocking function in our traversal, we simply wrap it in our aggregate and return the sub-tree as an MSE. For example, the expression $\sum_{j}\sqrt{A_{ij}B_{jk}}$ cannot be rewritten as $\sqrt{\sum_{j}A_{ij}\sum_{j}B_{jk}}$ because $\sqrt{}$ is a blocking function.
@@
Discussion. Galley builds upon and extends the FAQ framework for optimizing conjunctive queries with aggregation(Khamis et al., 2016). This framework explored the optimization of queries with the following form, where each $\bigoplus^{(i)}$ is either equal to or forms a semi-ring with $\otimes$:
@@
Though this framework captures many important problems, it lacks the flexibility needed to support a general tensor processing system. Consider a slightly modified version of the SDDMM kernel:
$\sum_{j}A_{ik}(B_{ij}+C_{jk})$
This expression is not an FAQ query because it mixes addition and multiplication in the pointwise expression. Similarly, our logistic regression example, $\sigma(\sum_{j}X_{ij}\theta_{j})$, cannot be expressed as an FAQ because the aggregate occurs within $\sigma$. Galley extends this framework by accommodating arbitrary pointwise function composition and arbitrary placement of aggregates within expressions.
@@ 5.5. Restricted Elimination Orders
Before we can search for the optimal elimination order, we must define the space of valid elimination orders. Depending on the input’s structure, the order in which indices can be eliminated might be restricted. This could be due to non-commutative aggregates or aggregate placement. The former occurs when an aggregate wraps another aggregate it does not commute with. For example, if the expression is $\max_{i}\sum_{j}A_{ij}$, we must perform the summation before handling the maximum because max and $\sum$ do not commute. The placement issue arises when an aggregate wraps another aggregate but cannot reach it via the traversal described above, e.g., $\sum_{i}\sqrt{\sum_{j}A_{ij}}$; in this case, the inner aggregate must be performed first. Collectively, these restrictions form a partial ordering on the index variables that must be respected when we enumerate elimination orders.
5.6. Search Algorithms
@@
With the VE approach, we have simplified the complicated issue of high-level optimization to the discrete problem of choosing an optimal order on the aggregated index variables. We start by revisiting our example from Fig. 5. The input query is the following,
    Query(X, Map($\sigma$,
                Agg(+,p,c,j,
                    Map(*,S[i,p,c],
                        Map(+,
                            Map(*, P[p,j], $\theta$[j]),
                            Map(*, C[c,j], $\theta$[j]))))))
The elimination order for this expression is an ordering of the indices $\{p,c,j^{1},j^{2}\}$. Note that $j^{1}$ and $j^{2}$ are virtual indices, introduced to represent the two paths that $j$ can be pushed down in the AET due to the addition operator. Galley’s logical optimizer searches through these possible orders to find the most efficient one. In this case, it would choose $[j^{1},j^{2},p,c]$, resulting in the following logical plan,
    Query(A1, Agg(+, j, Map(*, P[p,j], $\theta$[j])))
    Query(A2, Agg(+, j, Map(*, C[c,j], $\theta$[j])))
    Query(A3, Agg(+, p, c, Map(*, S[i,p,c],
                                Map(+, A1[p], A2[c]))))
    Query(X, Map($\sigma$, A3[i]))
We now present two algorithms to search for that optimal order using the tools described above.
@@
Greedy. The greedy approach simply chooses the cheapest index to aggregate at each point by finding the minimal sub-query for each index and computing its cost. The cheapest index’s minimal sub-query is removed from the expression and appended to the logical plan; it is then replaced in the remaining query with an alias to the result. This continues until no aggregates remain in the expression.
@@
Branch-and-Bound. The branch-and-bound approach computes the optimal variable order and occurs in two steps. The first step uses the greedy algorithm to produce an upper bound on the cost of the overall plan; the second performs a dynamic programming algorithm. In the dynamic programming step, the keys of the memo table are unordered sets of indices, and the values are a tuple of partial elimination orders, residual queries, and costs. The algorithm initializes the table with the empty set and a cost of zero. At each step, it iterates through table entries and attempts to aggregate out another index. It then uses the cost bound from the first step to prune entries from the memo table whose cost exceeds the bound; doing so is valid because costs monotonically increase as more indices are added to the set. At the end of this step, the algorithm returns the index order associated with the full set of indices.
@@ 6. Physical Optimizer
Each query in the logical dialect roughly corresponds to a single loop nest and materialized intermediate. However, several decisions remain about how the kernel is computed, including: (1) the loop order over the indices, (2) the format (i.e., layout) of the result, and (3) the protocol (i.e., search algorithm) for accessing each index of each input. The physical optimizer makes these decisions.
@@ 6.1. Loop Order
The loop order of a tensor kernel determines the order of access to the inputs. An optimal order results in early pruning of iterations due to early intersection of sparse inputs. Intuitively, this is similar to selecting a variable order for a worst-case optimal join algorithm. To find the optimal order, Galley has a cost model for loop orders based on sparsity estimates and an optimization algorithm that uses this cost model to guide its search.
@@
The cost of a loop order equals the number of iterations each level of the loop nest incurs.
@@
Consider matrix chain multiplication over three sparse matrices, $A$, $B$, and $C$, where
@@
Suppose that $A$ has only a single non-zero entry and that $B$ and $C$ have 5 non-zero entries per column and per row. In this case, the loop order $ijkl$ is significantly more efficient than $lkji$. In the former, the first two loops, over $i$ and $j$, incur only a single iteration because they are bounded by the size of $A$. The third and fourth incur $5$ and $5^{2}$ iterations, respectively, because there are only $5$ non-zero $k$’s per $j$ in $B$ and $5$ non-zero $l$’s per $k$ in $C$. In the latter, the first two loops iterate over the full matrix $C$ despite most of those iterations not leading to useful computation.
The last piece of our cost model is the cost of transposition. If an input’s index order is not concordant with the loop order, it must be transposed before the query can be executed. This imposes a cost that is linear in that input’s size. We ignore the cost of transposing intermediates because we they are always materialized in a concordant order.
Formally, let $Q$ be the pointwise expression in our kernel, and let $Q_{(i_{1},\ldots,i_{k})}$ be the restriction of that expression to just the sequence of index variables $i_{1},\ldots,i_{k}$. Further, let $\mathbf{A}^{(i_{1},\ldots,i_{k})}$ be the set of input tensors which are not concordant with $i_{1},\ldots,i_{k}$. Then, we can define the cost of a loop order as follows,
$cost(Q,(i_{1},\ldots,i_{k}))\approx\sum_{j=1}^{k}nnz(Q^{(i_{1},\ldots,i_{j})})+\sum_{A\in\mathbf{A}^{(i_{1},\ldots,i_{k})}}|A|$
In practice, we further refine this model to take into account the number and kind of tensor accesses at each level of the loop nest.
Optimization Algorithm.
To optimize the loop order, we combine this cost model with a branch-and-bound, dynamic programming algorithm. In the first pass, the optimization algorithm selects the cheapest loop index at each step until reaching a full loop order. This produces an upper bound on the optimal execution cost, which the algorithm uses to prune loop orders in the second step. This step applies a dynamic programming algorithm. Taking inspiration from Selinger’s algorithm for join ordering, each key in the DP table is a set of index variables and a set of inputs. The former represent the loops that have been iterated so far, and the latter represents a set of inputs that must be transposed.
@@ 6.2. Intermediate Formats
Once the loop order has been determined, the physical optimizer selects the optimal format for each query’s output. First, Galley selects the order of the indices in the output to be concordant with either the loop order of the kernel where it will be consumed or the order requested by the user. It then selects a format for each index of the output (e.g., dense vector, hash table, etc.). Two factors affect this decision: (1) the kind of writes being performed (sequential vs random) and (2) the sparsity of the tensor at this index. The former is important because many formats (e.g., sorted list formats) permit only sequential construction. These formats can only be used if the output indices up to that point form a prefix of the loop order.
When considering sparsity, Galley balances the fact that denser formats tend to be more efficient due to their memory locality and simplicity, while sparser formats are asymptotically smaller for sparse outputs. To describe this trade-off, we hand selected sparsity cutoffs between fully sparse, bytemap, and fully dense formats. To determine a particular output index’s format, the physical optimizer first determines the sparsity at this index level and uses our cutoffs to determine which category of formats to consider. Then, it checks whether sequential or random writes are being performed and selects the most efficient format that supports the write pattern.
@@ 6.3. Merge Algorithms
The final decision the physical optimizer makes concerns the algorithm it will use to perform each loop’s intersection. While there are more complex strategies, we adopt instead a minimal approach and select a single input to iterate over for each loop. The physical optimizer then probes into the remaining inputs. It makes this selection by estimating the number of non-zero indices that each input has, conditioned on the indices in the outer loops. This resembles the approach taken in (Wang et al., 2024) for optimizing WCOJ.
6.4. Common Sub-Expression Elimination
Galley takes a straightforward approach to avoiding redundant computation. Once a physical plan has been generated, the right hand side of each physical query is canonicalized and hashed. When two physical queries result in the same hash, the latter query is removed from the plan and all references to it are replaced with a reference to the result of the former. This is helpful for caching small computations like transpositions, but it is also useful for reducing the overhead of applying distributivity which often results in duplicate sub-expressions.
@@ 7. Sparsity Estimation
We now describe how Galley performs the sparsity estimation that guides our logical and physical optimizers. First, we explore the subtle correspondence between sparsity and cardinality estimation. We then present a minimal interface for sparsity estimation inspired by this correspondence, after which we examine two implementations of this framework, i.e., the uniform estimator and the chain bound.
7.1. Sparsity and Cardinality Estimation
Sparsity estimation is highly related to cardinality estimation in databases. However, translating methods for the latter to the former requires analyzing the algebraic properties of our tensor programs. For example, let $A_{ij}$ and $B_{jk}$ be sparse matrices with a fill value of 00, and let $R_{A}(I,J)$ and $R_{B}(J,K)$ be relations that store the indices of their non-zero entries. Assume we are performing the following,
@@ 7.1. Sparsity and Cardinality Estimation
In this case, the number of non-zero values in $C$ is precisely equal to the size of the conjunctive query
@@
The correspondence results from the fact that 00 is the annihilator of multiplication (i.e., $x*0=0\forall x$), so any non-zero entry $ijk$ in the output must correspond to a non-zero $ij$ in $A$ and a non-zero $jk$ in $B$. Consider the following instead:
@@
In this case, a nonzero $ijk$ in the output can result from a non-zero $ij$ in $A$ or a non-zero $jk$ in $B$. In traditional relational algebra, where relations are over infinite domains, this kind of disjunction would result in an infinite relation. However, tensors have finite dimensions, so we can introduce relations that represent the finite domains of each index, e.g., $D_{i}=\{1,...,n_{i}\}$. This lets us represent the index relation of the output as
@@
Finally, we can translate aggregations to the tensor setting as projection operations. Given the statement
@@
we can express the non-zeros entries of $C$ as
@@ 7.2. The Sparsity Statistics Interface
We use our statistics interface to annotate an expression with stats objects at every node of the AST in a bottom-up fashion. Each stats object then represents the sparsity patterns of the intermediate tensor output from that node. Surprisingly, to support sparsity estimation over the varied workloads and arbitrary operators of tensor algebra, we need to implement only a few core functions: (1) a constructor, which produces statistics from a materialized tensor for Input and Alias nodes, (2) a function for annihilating Map nodes (i.e., those whose children’s fill values are the annihilator of its pointwise function), which merges the children’s statistics, (3) a function for non-annihilating Map nodes, which merges the children’s statistics, (4) a function for Agg, which adjusts the input’s statistics to reflect an aggregation over some set of indices, and (5) an estimation procedure, which estimates the number of non-fill entries based on statistics about a tensor.